OpenStudy (anonymous):

Identify the 42nd term of an arithmetic sequence where a1 = -12 and a27 = 66. Question 6 options: 1) 70 2) 72 3) 111 4) 114

OpenStudy (angela210793):

a27=a1+26d 66=-12+26d 26d=78 d=3 a42=a1+41*3=-12+123=111